OpenStudy (anonymous):

how do i find the domain of the function: F(x)=square root of x^2-4/ x^2+x-6. I know im suppose to cancel out both x^2 but i got stuck on the next step.

hartnn (hartnn):

hi fairy princess :) is the function \(F(x)=\dfrac{\sqrt{x^2-4}}{x^2+x-6}\) ?

OpenStudy (anonymous):

hey and yes thats how the problem should look like, my bad

hartnn (hartnn):

no problem :) do you know how to get the domain ?

hartnn (hartnn):

of any function..

OpenStudy (anonymous):

eh....to be honest its confusing. could u give me a example of how to find the domain? maybe that would help

hartnn (hartnn):

sure! if f(x) = 1/(x+2) the denominator cannot be = 0 ! so, x\(\ne-2\) and -2 is excluded from the domain. for f(x) = \(\sqrt{x+2}\) the quantity under square root sign must be non-negative ! (so that x is real number) so, \(x+2\ge 0 \implies x\ge -2\) got these 2 examples ???

hartnn (hartnn):

so that f(x) is real number ***

OpenStudy (anonymous):

ohh...yeah i think i get it. like if its negative, it would make it undefined right?

hartnn (hartnn):

yup, undefined or complex(to be precise) so lets take your F(x) the denominator cannot be 0, right ? so can you solve for x^2+x-6=0 ?? quadratic equations?

OpenStudy (anonymous):

yeah that way we'll find a answer to x

hartnn (hartnn):

the 2 values you get for 'x' will be excluded from the domain, so, can you find those 2 values of x ?

OpenStudy (anonymous):

ohhh like x= positive or negatvie answer?

OpenStudy (anonymous):

but i thought x^2 over the other x^2 would cancel both out each other.

hartnn (hartnn):

not actually... \(\Large \dfrac{x^2+stuff}{x^2+otherstuff}\ne \dfrac{stuff}{otherstuff}\)

OpenStudy (anonymous):

oh okay thats where i was getting confused at

hartnn (hartnn):

cool, then what about x^2+x-6 = 0 ?? can u solve it ? find 2 numbers whose sum is +1 and product = - 6 ???

OpenStudy (anonymous):

wouldnt that be -7+1?

hartnn (hartnn):

no, i need sum to be 1 and product to be -6 try numbers like 2,3,-2,-3 ...

OpenStudy (anonymous):

-3?

hartnn (hartnn):

2 numbers ?

OpenStudy (anonymous):

-2 and 3

hartnn (hartnn):

correct!

hartnn (hartnn):

so, x^2+x-6 = 0 will be (x+3)(x-2) = 0 got this ?

OpenStudy (anonymous):

yay! okay so basically its finding the sum of 1 and that'll help get negative 6. and yeah i got it

hartnn (hartnn):

cool, so (x+3)(x-2) = 0 gives x=-3, x=2 these values will be excluded from the domain, because they will make the denominator =0 got this part?

OpenStudy (anonymous):

yeah and its cause the denominator cant be zero?

hartnn (hartnn):

right! denominator = 0 is undefined

hartnn (hartnn):

now for numerator, which is under square root the quantity under square root must be non-negative, right? so, solve for \(x^2-4 \ge 0\) can u ?

OpenStudy (anonymous):

right and yeah all u gotta do is add four then \[x ^{2}=4\]

OpenStudy (anonymous):

then youll square root both and youll get postive or negative two right?

hartnn (hartnn):

good! so, \(x\ge 2, \: x\le-2\) any doubts ?

OpenStudy (anonymous):

hmmmm no not really, i get it now much better. just my teacher didnt put all of that into two parts. like doing the denominator and the numerator

hartnn (hartnn):

ok, and lastly we write everything together in the domain \(domain \rightarrow (x\ge 2,x\le -2, x\ne 2, x\ne -3)\rightarrow (x> 2,x\le -2, x\ne -3)\)

hartnn (hartnn):

and since you are new here, \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{P}}\color{purple}{\text{E}}\color{red}{\text{N}}\color{orange}{\text{S}}\color{#e6e600}{\text{T}}\color{green}{\text{U}}\color{blue}{\text{D}}\color{purple}{\text{Y}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{}}\end{array} \] have fun learning :)

OpenStudy (anonymous):

thank you for the help and yeah ill try to have fun learning xD Thanks for the welcoming also

hartnn (hartnn):

most welcome ^_^